HELIUM PRESSURE BROADENING OF HDO BETWEEN 2 AND 80 K

\underline{T. J. RONNINGEN} and F. C. DE LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210.

Below 80 K the pressure broadening cross section for HDO-He collisions is predicted to be less than 10 Å^{2}. This is quite small for a microwave transition and suggests that the line may, at low pressures, narrow with increasing pressure (Dicke narrowing). We have employed collisional cooling to measure the lineshape of the HDO 1_{01} − 1_{11} transition between 2 and 80 K. Initial results for the high pressure data indicate that the theoretical predictions for the cross section are valid. A more detailed analysis that investigates non-Lorentzian contributions to the lineshape will be discussed.